A

surface of the earth. When the satellite orbits around the earth at a height equal to the diameter of the earth (above the surface of the earth), the magnitude of the gravitational force on the satellite will be:

satellite experiences a gravitational force with magnitude F when it is on the

A 3F

B 2F

C 1/3 F

D 1/9 F


1
Expert's answer
2021-11-29T11:45:56-0500

The height equal to Earth's diameter (or "2R") above the surface of the Earth is actually equivalent to the distance of "3R" from the center of the planet. As we know, the gravitational force (at the surface) is


"F=G\frac{Mm}{R^2}."

At distance of "3R":


"F_1=G\frac{Mm}{(3R)^2}=\frac F9."